Find four consecutive odd integers if the product of the two smaller integers is 112 less than the product of the two larger integers. How do I set this up?
Answer by scott8148(6628)

if n is the 1st integer, then the other three are n+2, n+4, and n+6

n (n+2) = (n+4) (n+6) - 112

n^2 + 2n = n^2 + 10n - 88

88 = 8n ___ 11 = n

11, 13, 15, 17